# A Survey of Reinforcement Learning For Economics

### 23 Mar 2026 v4

@@ A Survey of Reinforcement Learning For Economics
This survey (re)introduces reinforcement learning methods to economists. The curse of dimensionality limits how far exact dynamic programming can be effectively applied, forcing us to rely on suitably “small” problems or our ability to convert “big” problems into smaller ones. While this reduction has been sufficient for many classical applications, a growing class of economic models resists such reduction. Reinforcement learning algorithms offer a natural, sample-based extension of dynamic programming, extending tractability to problems with high-dimensional states, continuous actions, and strategic interactions. I review the theory connecting classical planning to modern learning algorithms and demonstrate their mechanics through simulated examples in pricing, inventory control, strategic games, and preference elicitation. I also examine the practical vulnerabilities of these algorithms, noting their brittleness, sample inefficiency, sensitivity to hyperparameters, and the absence of global convergence guarantees outside of tabular settings. The successes of reinforcement learning remain strictly bounded by these constraints, as well as a reliance on accurate simulators. That said, when guided by economic structure, reinforcement learning provides a flexible and innovative framework. It stands as an imperfect, but promising, addition to the computational economist’s toolkit. A companion survey (Rust and Rawat, 2026b) covers the inverse problem of inferring preferences from observed behavior. All simulation code is publicly available. [^1] https://github.com/rawatpranjal/survey-of-reinforcement-learning-in-economics
@@ 1 Introduction
This survey (re)introduces reinforcement learning to economists. I review the theoretical connections between dynamic programming and reinforcement learning, demonstrating how value iteration, Q-learning, and policy gradient methods are common solution methods to the same class of optimization problems. I then examine applications across several domains, including control problems, pricing and inventory management; structural economic models with high-dimensional state spaces; strategic games in which multi-agent algorithms compute equilibria under imperfect information; bandit problems in which economic structure yields tighter regret bounds; and preference learning. The exposition combines formal theory, practical applications and computational illustration.
@@
This survey focuses on less-surveyed intersections between reinforcement learning and economics, including the shared theoretical foundations, structural estimation, strategic interaction, bandit problems with economic structure, preference learning, and causal inference. Algorithmic collusion, in which independent pricing algorithms learn to sustain supra-competitive prices (Calvano et al., 2020), is treated in a companion thesis chapter (Rawat, 2026) and omitted here. Reinforcement learning and deep learning methods for solving macroeconomic models with heterogeneous agents constitute a growing literature with dedicated methodological treatments (Atashbar and Shi, 2022), (Maliar et al., 2021), and (Fernández-Villaverde et al., 2024). Portfolio optimization, optimal execution, and asset pricing via reinforcement learning form a large body of work surveyed comprehensively elsewhere (Hambly et al., 2023). The inverse problem of inferring preferences from observed behavior using inverse reinforcement learning and structural estimation is treated in a companion survey (Rust and Rawat, 2026).
The survey addresses the forward problem, that is, computing optimal policies given a known or simulated environment. Chapter 1 traces the parallel historical development of dynamic programming and reinforcement learning. Chapter 2 develops unified theory connecting planning and learning. Chapters 3 and 4 apply reinforcement learning to control problems and economic models. Chapter 5 examines strategic games. Chapter 6 addresses bandit problems. Chapter 7 discusses reinforcement learning from human feedback. Chapter 8 connects reinforcement learning to causal inference. Chapter 9 concludes.
@@ 2 Two Cultures of Sequential Decision-Making
Economics and reinforcement learning both study sequential decision-making under uncertainty, but they descend from different intellectual traditions. Economics is fundamentally an inference culture. Its central task is to understand the world. An economist’s “model” is a specification of preferences, beliefs, constraints, and an equilibrium concept. The RL tradition is instead a control culture. Its central task is to act in the world. An RL researcher’s “model” is a transition kernel $P(s^{\prime}|s,a)$ and a reward function $r(s,a)$. These are different mathematical objects serving fundamentally different scientific purposes.
The two cultures maintain different relationships with data. Economists work primarily with observational data, and endogeneity is the central obstacle. Identification is its defining challenge which involves dealing with functional forms, equilibrium definitions, instrumental variables, regression discontinuities, and natural experiments. RL researchers have traditionally enjoyed what might be called simulator omnipotence. They often own the data-generating process and can create the variation in their data and fit any manner of nonparametric functions through millions of simulated samples. They face computational rather than statistical constraints. This difference shapes the cultures, from algorithm design to what counts as a valid result and what is the standard of evidence.
The two fields also differ in their treatment of human agency and in the meaning of basic terms like “agent,” “learning,” “model,” and “policy,” as detailed in the subsections that follow.
The two cultures have a closely related and shared vocabulary but with divergent meanings. And this creates persistent confusion for researchers crossing from one field to the other. This chapter attempts to provide a systematic translation and notes common sources of confusion.
@@ 2.1 Core Distinctions
RL organizes its core problems along axes that do not map directly onto well known categories. In RL, Prediction refers to estimating $V^{\pi}(s)$ or $Q^{\pi}(s,a)$ for a fixed policy $\pi$ (policy evaluation), not forecasting observable variables. Control refers to finding the policy $\pi^{*}$ that maximizes expected discounted return (policy optimization), not the inclusion of regressors. Because prediction concerns evaluating a specific policy, a closely related question is whether the data was generated by the policy being studied. On-policy methods evaluate and improve the same policy that generates the data. Off-policy methods learn about a target policy $\pi$ from data generated by a different behavioral policy $\mu$. This distinction is central to causal inference (Section 12), where off-policy evaluation is precisely counterfactual policy evaluation.222“Counterfactual” here is used in the interventional sense, asking “what would happen if we deployed policy $\pi$ instead of $\mu$?” This is distinct from the structural counterfactual in Oberst and Sontag (2019), which conditions on the specific realized trajectory and asks what would have happened to this individual under a different action, requiring a fully specified structural causal model rather than just the observational distribution under $\mu$.
@@
A separate question concerns how the algorithm accesses experience. Online RL learns while interacting with the environment, collecting new data as a consequence of its own actions. Offline RL (also called batch RL) learns exclusively from a fixed dataset of previously collected transitions, with no ability to gather additional samples. The offline setting is closer to standard empirical work, where the dataset is given and the analyst cannot run new experiments; the online setting corresponds more closely to adaptive experimental design or sequential decision problems. Note that “online” in RL carries no timing constraint; it means only that the agent generates fresh experience. Real-time RL, by contrast, imposes hard deadlines on the perception-action loop, as in robotics or mechatronics. Every real-time RL system is online, but most online RL (games, recommender systems) are not real-time.
@@
Aside from data access is the question of what the algorithm does with that data. In RL, the word model refers strictly to the environment’s dynamics, the transition kernel $P(s^{\prime}|s,a)$ and reward function $r(s,a)$. A model-based RL algorithm explicitly constructs or is given a mathematical representation of $P$ and $r$, then computes a policy by planning through that representation (for example, using a simulator or the known rules of a game, as in AlphaZero). A model-free algorithm computes the value function or policy directly from experienced transitions without ever building an explicit representation of the transition probabilities. Crucially, “model-free” need not mean the algorithm lacks complete access to any model of the environment. A model-free algorithm could interact with a simulator that internally implements a complete computerized model of the environment; the distinction is that the algorithm never extracts or plans through the transition probabilities, treating the simulator as a black box that merely returns sample transitions. However, it is possible that a “model-free” algorithm could also be implemented “in-field” where there is genuinely no access to a “model” of the environment but only direct access to the environment itself (for reasons discussed later, this is rarely done).
One might believe that an analog to “model-free” RL algorithms might be reduced-form or a statistical model that only requires a few assumptions (exogeneity, etc.) to make inference on parameters of interest. But this is a false analogy, as regardless of “model-free” or “model-based” we are talking about learning optimal policies within a MDP (this itself demands a certain structure on the problem including the Markovian assumption that the state contains all relevant information). One could believe that “model-based” algorithms are analogous to structural estimation, which involves the specification of preferences, technology, beliefs, and an equilibrium concept, all of which give rise to specific transition dynamics and reward functions. This is also a false analogy. In RL terminology, for an algorithm to be “model-based”, the conditions are very precise, all we need is a given representation of $P$ and $r$ (regardless of whether you do or don’t make any structural economic assumptions about human or firm behavior or their interactions). All this to say, in RL, the model generally only refers to transition matrices $P$ and rewards $r$ within an MDP; whereas in economics it refers to a set of agents, preferences, exclusion restrictions, and equilibrium concepts. Therefore, it is possible to do structural estimation of an “economic model” with a lot of structure but by using a “model-free” RL algorithm as a computational tool to find the equilibrium, as in Section 8.
@@
Every RL system passes through two phases. In the training phase, the agent interacts with an environment (simulated or real) and updates its parameters. In the execution phase (also called the deployment phase), the policy is frozen and used to make decisions without further updates. This distinction is critical for interpreting “online.” Online training in RL almost always takes place inside a simulator (and not in the “real world”). AlphaGo Zero trained online through millions of self-play games; when it faced Lee Sedol, its weights were frozen and it was purely executing its trained policy. Some deployed systems in Section 7 followed this pattern cleanly; DiDi’s dispatch system trained a value function from historical trip data, then deployed with fixed weights. Others blur the boundary. The hotel revenue management system in Section 7.2 updated Q-values from realized returns after each completed episode during live operation, making it an in-field learner rather than a frozen executor.333“In-field” is not standard RL vocabulary. We introduce it here to distinguish live-market online learning, where exploration has real economic cost, from the far more common case of online learning inside a simulator. Bandit algorithms (Section 10) also learn in-field by design, updating demand estimates from real customer responses during deployment.
@@
A related terminological hazard concerns the word inference. In machine learning, “inference” refers to executing a frozen model, a forward pass producing outputs from inputs. In economics, “inference” means statistical inference, the construction of standard errors, confidence intervals, and hypothesis tests. This survey uses “inference” exclusively in the econometric sense and “execution” or “deployment” when referring to applying a trained model (whether in a computer or in field). Established terms such as “Bayesian inference” and “variational inference,” which refer to inference about parameters or distributions, are used where appropriate. The key takeaway is that most RL convergence results and sample complexity guarantees refer to the training phase, and interpreting them as claims about deployed performance requires additional argument. Table 1 summarizes these distinctions.
@@
Table 1: The reinforcement learning lifecycle grid. Most RL research operates in the top-left cell. When economists hear “online,” they typically picture the bottom-left.
@@
A typical economic RL pipeline moves through the grid sequentially, from pre-training on historical logs (middle-left) to refinement in a simulator (top-left) to deployment with frozen weights (bottom-right). Bandits illustrate this fluidity; even a bandit algorithm that will ultimately learn in-field is typically calibrated in simulation and tuned on historical logs before any live deployment, because in-field exploration incurs real financial cost. The systems that do operate in-field arrive with exploration parameters, initial policies, and demand priors shaped by extensive offline preparation.
@@ 2.2 Overlapping Terminology
The word agent refers to different entities. In economics, an agent is always a human decision-maker (a consumer, worker, or firm) or a social planner whose choices are the object of study. In RL, “the agent” is the learning algorithm itself, or more precisely an algorithm deployed on behalf of a human decision-maker, and the human, if present at all, is part of the environment providing reward signals.
@@
In RL the environment is a formal object encompassing everything outside the agent (the DGP, other agents, market clearing conditions), whereas in economics “environment” refers more loosely to market structure or institutional rules. [^4] A source of confusion is generative model. In machine learning broadly, this means a model of the joint distribution $P(X,Y)$ or a synthetic data generator (GANs, diffusion models). In RL theory, a generative model is a simulator oracle that, given any $(s,a)$, returns a sample $s^{\prime}\sim P(\cdot|s,a)$ and reward $r(s,a)$; the usage implies random-access simulation, stronger than sequential online interaction but weaker than knowing $P$ analytically.. RL speaks of rewards where economists speak of utility. The mapping is not exact: a reward $r(s,a)$ is a known, externally specified function that the algorithm maximizes, whereas utility $u(x,d)$ in economics is a primitive of preferences that the analyst must recover or estimate from observed choices.
@@
In RL, the return $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}$ is the discounted sum of future rewards from time $t$ onward, the random variable whose expectation defines the value function. The RL usage is closer to what economists call the “present discounted value” of a stream of payoffs. A single complete sequence of interactions from an initial state to termination, $(s_{0},a_{0},r_{1},s_{1},\ldots,s_{T})$, is called an episode (synonymously, trajectory or rollout). [^5] The closest econometric analogs are a single panel unit’s time series, a realization of a stochastic process, or one “history” in a dynamic model. Where an econometrician speaks of the outcome, meaning the dependent variable $Y$ in a regression, RL has no single analog: the reward $r_{t}$ is the per-period outcome, the return $G_{t}$ is the cumulative outcome, and the value function $V^{\pi}(s)$ is the expected cumulative outcome conditional on state.
@@
Learning carries many distinct meanings in the two fields. In economic theory (Bayesian learning, adaptive expectations), learning refers to agents forming and refining beliefs about unknown parameters of their environment. In supervised machine learning, learning means statistical estimation, fitting the weights of a parameterized model to minimize a loss function over data. In reinforcement learning, “learning” is primarily computation. When an RL agent “learns” a Q-function, it is executing a recursive stochastic approximation algorithm to find the fixed point of the Bellman operator. We say the algorithm is “learning” because it improves its policy iteratively through simulated or real experience, but mathematically it is solving a fixed-point problem. In many applications throughout this survey, particularly Section 8, the RL algorithm is simply a numerical method for solving the Bellman equation; no actual human-like learning from experience is taking place. Finally, while RL draws inspiration from animal psychology (Section 3.1), it is a drastic simplification of biological learning. Tabula rasa RL algorithms require millions of iterations of trial and error to discover policies that animals acquire rapidly. Real-world animal learning relies on innate priors, basic physical knowledge, and parental nurturing and should be distinct from RL-style “learning”.
@@
Macroeconomic adaptive learning has used ideas formally analogous to reinforcement learning for decades, though under different names and with different motivations. In the adaptive learning literature initiated by Marcet and Sargent (1989), boundedly rational agents update their beliefs about equilibrium parameters using recursive least-squares and other Robbins-Monro-type stochastic approximation algorithms. The convergence criterion in that literature, E-stability, evaluates whether the ordinary differential equation (ODE) associated with the mapping from the perceived to the actual law of motion is locally asymptotically stable at the rational expectations equilibrium. This fixed-point stability requirement is analogous to the contraction conditions governing the convergence of temporal-difference and Q-learning algorithms in RL. Borkar and Meyn (2000) make this mathematical connection explicit: they prove the convergence of Q-learning and actor-critic methods using the ODE method, explicitly citing Sargent (1993) as a parallel application of the exact same framework to boundedly rational economic agents.
@@
Active learning appears in both fields but means different things. In economics, active learning denotes Bayesian experimentation where an agent endogenously chooses what information to acquire, trading off the cost of exploration against the option value of better future decisions. In machine learning, active learning is a supervised learning protocol in which the algorithm selects which unlabeled examples to query an oracle for labels, minimizing annotation cost rather than maximizing cumulative reward. The term exploration is ubiquitous in RL but rarely used in economics. In RL, exploration is a modular subcomponent of a larger algorithm, a procedure for choosing informative actions that may be quite crude (such as $\varepsilon$-greedy random action selection) or more principled (UCB, Thompson sampling). The closest economic analog is optimal experimentation in the Bayesian bandit tradition (Rothschild 1974), where the value of information is derived endogenously from the agent’s dynamic program, not bolted on as a separate heuristic. Economists also use Bayesian learning to describe the same phenomenon, but always within a fully specified model of beliefs and preferences; in RL, exploration can be a purely algorithmic device with no decision-theoretic foundation.
@@
Bootstrapping in econometrics refers to Efron’s resampling method (Efron, 1979); in RL, it means updating a value estimate using another value estimate rather than a complete realized return, as when a TD algorithm uses the target $r_{t+1}+\gamma V(s_{t+1})$ that depends on the current, uncertain estimate $V(s_{t+1})$ (Sutton, 1988). Function approximation in RL refers to representing value functions or policies using parameterized function classes (linear combinations of basis functions, kernel methods, or neural networks), which econometricians will recognize as sieve estimation or nonparametric series estimation, the approximation of an unknown function by projection onto a finite-dimensional basis. Calibration carries unrelated meanings. In economics, calibration means choosing model parameters to match a set of empirical moments or stylized facts (Kydland and Prescott 1982). In machine learning, calibration refers to probability calibration, ensuring that predicted probabilities match observed frequencies, a statistical property of a classifier’s outputs.
@@
The term bandit itself carries different mathematical content across disciplines. The classical multi-armed bandit in statistics and economics (Thompson, 1933; Rothschild, 1974; Gittins, 1979) is a Bayesian sequential allocation problem. The state is the agent’s posterior belief over the unknown arm distributions, and the solution is the Gittins index, an optimal allocation rule derived from the theory of optimal stopping. In the RL and computer science literature, bandits are instead framed as frequentist regret-minimization problems. Algorithms such as UCB provide worst-case bounds on cumulative regret $\sum_{t=1}^{T}(\mu^{*}-\mu_{A_{t}})$ without requiring Bayesian priors. The two traditions ask fundamentally different questions, Bayesian optimality of the full sequential problem versus minimax regret rates over adversarial or stochastic environments, and their answers are not directly comparable. Section 10 adopts the regret framework because it connects more naturally to the sample complexity concerns that arise in economic field experiments.

### 24 Mar 2026 v5

@@ A Survey of Reinforcement Learning For Economics
This survey (re)introduces reinforcement learning methods to researchers in the social sciences. The curse of dimensionality limits how far exact dynamic programming can be effectively applied, forcing us to rely on suitably “small” problems or our ability to convert “big” problems into smaller ones. While this reduction has been sufficient for many classical applications, a growing class of economic models resists such reduction. Reinforcement learning algorithms offer a natural, sample-based extension of dynamic programming, extending tractability to problems with high-dimensional states, continuous actions, and strategic interactions. I review the theory connecting classical planning to modern learning algorithms and demonstrate their mechanics through simulated examples in pricing, inventory control, strategic games, and preference elicitation. I also examine the practical vulnerabilities of these algorithms, noting their brittleness, sample inefficiency, sensitivity to hyperparameters, and the absence of global convergence guarantees outside of tabular settings. The successes of reinforcement learning remain strictly bounded by these constraints, as well as a reliance on accurate simulators. That said, when guided by economic structure, reinforcement learning provides a flexible and innovative framework. It stands as an imperfect, but promising, addition to the researcher’s toolkit. A companion survey (Rust and Rawat, 2026b) covers the inverse problem of inferring preferences from observed behavior. All simulation code is publicly available. [^1] https://github.com/rawatpranjal/survey-of-reinforcement-learning-in-economics
@@ 1 Introduction
This survey (re)introduces reinforcement learning to researchers studying sequential decision problems. I review the theoretical connections between dynamic programming and reinforcement learning, demonstrating how value iteration, Q-learning, and policy gradient methods are common solution methods to the same class of optimization problems. I then examine applications across several domains, including control problems, pricing and inventory management; structural economic models with high-dimensional state spaces; strategic games in which multi-agent algorithms compute equilibria under imperfect information; bandit problems in which economic structure yields tighter regret bounds; and preference learning. The exposition combines formal theory, practical applications and computational illustration.
@@
This survey focuses on less-surveyed intersections between reinforcement learning and applied decision-making, including the shared theoretical foundations, structural estimation, strategic interaction, bandit problems with domain structure, preference learning, and causal inference. Algorithmic collusion, in which independent pricing algorithms learn to sustain supra-competitive prices (Calvano et al., 2020), is treated in a companion thesis chapter (Rawat, 2026) and omitted here. Reinforcement learning and deep learning methods for solving macroeconomic models with heterogeneous agents constitute a growing literature with dedicated methodological treatments (Atashbar and Shi, 2022), (Maliar et al., 2021), and (Fernández-Villaverde et al., 2024). Portfolio optimization, optimal execution, and asset pricing via reinforcement learning form a large body of work surveyed comprehensively elsewhere (Hambly et al., 2023). The inverse problem of inferring preferences from observed behavior using inverse reinforcement learning and structural estimation is treated in a companion survey (Rust and Rawat, 2026).
The survey addresses the forward problem, that is, computing optimal policies given a known or simulated environment. Chapter 1 traces the parallel historical development of dynamic programming and reinforcement learning. Chapter 2 develops unified theory connecting planning and learning. Chapters 3 and 4 apply reinforcement learning to control problems and applied models. Chapter 5 examines strategic games. Chapter 6 addresses bandit problems. Chapter 7 discusses reinforcement learning from human feedback. Chapter 8 connects reinforcement learning to causal inference. Chapter 9 concludes.
@@ 2 Two Cultures of Sequential Decision-Making
Two intellectual traditions both study sequential decision-making under uncertainty, but they descend from different intellectual traditions. The first is fundamentally an inference culture. Its central task is to understand the world. A “model” in this tradition is a specification of preferences, beliefs, constraints, and an equilibrium concept. The RL tradition is instead a control culture. Its central task is to act in the world. An RL researcher’s “model” is a transition kernel $P(s^{\prime}|s,a)$ and a reward function $r(s,a)$. These are different mathematical objects serving different scientific purposes.
The inference tradition concentrates its effort on specifying and estimating the objective function and the law of motion that governs the environment. The entire structural econometrics enterprise (demand estimation, dynamic discrete choice) is devoted to recovering these primitives from data, and the optimal policy is a byproduct that falls out once the primitives have been identified. The control tradition inverts this emphasis. Engineers typically take the objective and the dynamics as given (the cost function is specified, the plant physics are known or measurable) and focus on whether the optimal policy can be computed, approximated, and deployed under real-time and robustness constraints. The entire controls enterprise (PID, LQR, MPC, RL) is about computing and implementing the policy under different assumptions about what the agent knows and what it can compute.
The two cultures maintain different relationships with data. Econometricians work primarily with observational data, where endogeneity is the central obstacle; agents sort, markets clear, and unobservables correlate with the variables of interest. Identification, the question of whether the data can distinguish the true model from observationally equivalent alternatives, is the defining challenge, and it disciplines every modeling choice (functional forms, equilibrium definitions, instrumental variables, regression discontinuities, natural experiments). RL researchers have traditionally enjoyed what might be called simulator omnipotence. They own the data-generating process, can inject arbitrary variation through exploration policies, and can generate millions of on-policy rollouts at negligible marginal cost. Their binding constraint is computational (can the algorithm converge before the compute budget runs out?) rather than statistical (is the estimator consistent given the endogeneity structure of the data?). This asymmetry shapes everything downstream, from what counts as a valid result to what the word “model” means. To an econometrician, a model is a set of falsifiable restrictions on the joint distribution of observables and primitives. To an RL researcher, a model is a simulator you can call.
The two fields also share a vocabulary (“agent,” “learning,” “model,” “policy”) whose meanings diverge in ways that create persistent confusion. The subsections below provide a systematic translation.
@@ 2.1 Core Distinctions
In RL, prediction refers to estimating $V^{\pi}(s)$ or $Q^{\pi}(s,a)$ for a fixed policy $\pi$ (policy evaluation), not forecasting observable variables. Control refers to finding the policy $\pi^{*}$ that maximizes expected discounted return (policy optimization), not the inclusion of regressors. Because prediction concerns evaluating a specific policy, a closely related question is whether the data was generated by the policy being studied. On-policy methods evaluate and improve the same policy that generates the data. Off-policy methods learn about a target policy $\pi$ from data generated by a different behavioral policy $\mu$. This distinction is central to causal inference (Section 12), where off-policy evaluation is precisely counterfactual policy evaluation.222“Counterfactual” here is used in the interventional sense, asking “what would happen if we deployed policy $\pi$ instead of $\mu$?” This is distinct from the structural counterfactual in Oberst and Sontag (2019), which conditions on the specific realized trajectory and asks what would have happened to this individual under a different action, requiring a fully specified structural causal model rather than just the observational distribution under $\mu$.
@@
Online RL learns while interacting with the environment, collecting new data as a consequence of its own actions. Offline RL (also called batch RL) learns exclusively from a fixed dataset of previously collected transitions, with no ability to gather additional samples. The offline setting is closer to standard empirical work, where the dataset is given and the analyst cannot run new experiments; the online setting corresponds more closely to adaptive experimental design or sequential decision problems. Note that “online” in RL carries no timing constraint; it means only that the agent generates fresh experience. Real-time RL, by contrast, imposes hard deadlines on the perception-action loop, as in robotics or mechatronics. Every real-time RL system is online, but most online RL (games, recommender systems) are not real-time.
@@
In RL, the word model refers strictly to the environment’s dynamics, the transition kernel $P(s^{\prime}|s,a)$ and reward function $r(s,a)$. A model-based RL algorithm explicitly constructs or is given a mathematical representation of $P$ and $r$, then computes a policy by planning through that representation (for example, using a simulator or the known rules of a game, as in AlphaZero). A model-free algorithm computes the value function or policy directly from experienced transitions without ever building an explicit representation of the transition probabilities. “Model-free” need not mean the algorithm lacks access to any model of the environment. A model-free algorithm could interact with a simulator that internally implements a complete computerized model of the environment; the distinction is that the algorithm never extracts or plans through the transition probabilities, treating the simulator as a black box that merely returns sample transitions. However, it is possible that a “model-free” algorithm could also be implemented “in-field” where there is genuinely no access to a “model” of the environment but only direct access to the environment itself (for reasons discussed later, this is rarely done).
Neither “model-free” nor “model-based” maps onto the reduced-form versus structural distinction in econometrics. Both labels refer to whether the algorithm uses an explicit representation of $P$ and $r$, not to whether the analyst makes structural assumptions about preferences or equilibrium. A “model-based” RL algorithm needs only a representation of $P$ and $r$, regardless of whether those objects arise from structural assumptions about human behavior. Conversely, “model-free” does not mean “assumption-free”; both variants operate within an MDP, which itself imposes the Markov assumption on the state. In the inference tradition, “model” refers to a set of agents, preferences, exclusion restrictions, and equilibrium concepts. It is therefore possible to specify a rich structural model but solve for its equilibrium using a model-free RL algorithm as a computational tool, as in Section 8.
@@
Every RL system passes through two phases. In the training phase, the agent interacts with an environment (simulated or real) and updates its parameters. In the execution phase (also called the deployment phase), the policy is frozen and used to make decisions without further updates. This distinction is critical for interpreting “online.” Online training in RL almost always takes place inside a simulator (and not in the “real world”). AlphaGo Zero trained online through millions of self-play games; when it faced Lee Sedol, its weights were frozen and it was purely executing its trained policy. Some deployed systems in Section 7 followed this pattern cleanly; DiDi’s dispatch system trained a value function from historical trip data, then deployed with fixed weights. Others blur the boundary. The hotel revenue management system in Section 7.2 updated Q-values from realized returns after each completed episode during live operation, making it an in-field learner rather than a frozen executor.333“In-field” is not standard RL vocabulary. We introduce it here to distinguish live-market online learning, where exploration has real financial cost, from the far more common case of online learning inside a simulator. Bandit algorithms (Section 10) also learn in-field by design, updating demand estimates from real customer responses during deployment.
@@
The word inference is overloaded. In machine learning, “inference” refers to executing a frozen model, a forward pass producing outputs from inputs. In statistics, “inference” means statistical inference, the construction of standard errors, confidence intervals, and hypothesis tests. This survey uses “inference” exclusively in the statistical sense and “execution” or “deployment” when referring to applying a trained model (whether in a computer or in field). Established terms such as “Bayesian inference” and “variational inference,” which refer to inference about parameters or distributions, are used where appropriate. The key takeaway is that most RL convergence results and sample complexity guarantees refer to the training phase, and interpreting them as claims about deployed performance requires additional argument. Table 1 summarizes these distinctions.
@@
Table 1: The reinforcement learning lifecycle grid. Most RL research operates in the top-left cell. Readers from the inference tradition typically picture the bottom-left when they hear “online.”
@@
A typical applied RL pipeline moves through the grid sequentially, from pre-training on historical logs (middle-left) to refinement in a simulator (top-left) to deployment with frozen weights (bottom-right). Bandits illustrate this fluidity; even a bandit algorithm that will ultimately learn in-field is typically calibrated in simulation and tuned on historical logs before any live deployment, because in-field exploration incurs real financial cost. The systems that do operate in-field arrive with exploration parameters, initial policies, and demand priors shaped by extensive offline preparation.
@@ 2.2 Overlapping Terminology
In the inference tradition, an agent is always a human decision-maker (a consumer, worker, or firm) or a social planner whose choices are the object of study. In RL, “the agent” is the learning algorithm itself, or more precisely an algorithm deployed on behalf of a human decision-maker, and the human, if present at all, is part of the environment providing reward signals.
@@
In RL the environment is a formal object encompassing everything outside the agent (the DGP, other agents, market clearing conditions), whereas in the inference tradition “environment” refers more loosely to market structure or institutional rules. [^4] A source of confusion is generative model. In machine learning broadly, this means a model of the joint distribution $P(X,Y)$ or a synthetic data generator (GANs, diffusion models). In RL theory, a generative model is a simulator oracle that, given any $(s,a)$, returns a sample $s^{\prime}\sim P(\cdot|s,a)$ and reward $r(s,a)$; the usage implies random-access simulation, stronger than sequential online interaction but weaker than knowing $P$ analytically.. RL speaks of rewards where the other tradition speaks of utility. The mapping is not exact: a reward $r(s,a)$ is a known, externally specified function that the algorithm maximizes, whereas utility $u(x,d)$ in the inference tradition is a primitive of preferences that the analyst must recover or estimate from observed choices.
@@
In RL, the return $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}$ is the discounted sum of future rewards from time $t$ onward, the random variable whose expectation defines the value function. The RL usage is closer to what is called the “present discounted value” of a stream of payoffs. A single complete sequence of interactions from an initial state to termination, $(s_{0},a_{0},r_{1},s_{1},\ldots,s_{T})$, is called an episode (synonymously, trajectory or rollout). [^5] The closest statistical analogs are a single panel unit’s time series, a realization of a stochastic process, or one “history” in a dynamic model. Where a statistician speaks of the outcome, meaning the dependent variable $Y$ in a regression, RL has no single analog: the reward $r_{t}$ is the per-period outcome, the return $G_{t}$ is the cumulative outcome, and the value function $V^{\pi}(s)$ is the expected cumulative outcome conditional on state.
@@
Learning has several distinct meanings. In decision theory (Bayesian learning, adaptive expectations), learning refers to agents forming and refining beliefs about unknown parameters of their environment. In supervised machine learning, learning means statistical estimation, fitting the weights of a parameterized model to minimize a loss function over data. In reinforcement learning, “learning” is primarily computation. When an RL agent “learns” a Q-function, it is executing a recursive stochastic approximation algorithm to find the fixed point of the Bellman operator. We say the algorithm is “learning” because it improves its policy iteratively through simulated or real experience, but mathematically it is solving a fixed-point problem. In many applications throughout this survey, particularly Section 8, the RL algorithm is simply a numerical method for solving the Bellman equation; no actual human-like learning from experience is taking place. Finally, while RL draws inspiration from animal psychology (Section 3.1), it is a drastic simplification of biological learning. Tabula rasa RL algorithms require millions of iterations of trial and error to discover policies that animals acquire rapidly. Real-world animal learning relies on innate priors, basic physical knowledge, and parental nurturing and should be distinct from RL-style “learning”.
@@
Adaptive learning in macroeconomics has used ideas formally analogous to reinforcement learning for decades, though under different names and with different motivations. In the adaptive learning literature initiated by Marcet and Sargent (1989), boundedly rational agents update their beliefs about equilibrium parameters using recursive least-squares and other Robbins-Monro-type stochastic approximation algorithms. The convergence criterion in that literature, E-stability, evaluates whether the ordinary differential equation (ODE) associated with the mapping from the perceived to the actual law of motion is locally asymptotically stable at the rational expectations equilibrium. This fixed-point stability requirement is analogous to the contraction conditions governing the convergence of temporal-difference and Q-learning algorithms in RL. Borkar and Meyn (2000) make this mathematical connection explicit: they prove the convergence of Q-learning and actor-critic methods using the ODE method, explicitly citing Sargent (1993) as a parallel application of the exact same framework to boundedly rational agents.
@@
Active learning appears in both fields but means different things. In the inference tradition, active learning denotes Bayesian experimentation where an agent endogenously chooses what information to acquire, trading off the cost of exploration against the option value of better future decisions. In machine learning, active learning is a supervised learning protocol in which the algorithm selects which unlabeled examples to query an oracle for labels, minimizing annotation cost rather than maximizing cumulative reward. The term exploration is ubiquitous in RL but rarely used in the inference tradition. In RL, exploration is a modular subcomponent of a larger algorithm, a procedure for choosing informative actions that may be quite crude (such as $\varepsilon$-greedy random action selection) or more principled (UCB, Thompson sampling). The closest analog in the other tradition is optimal experimentation in the Bayesian bandit tradition (Rothschild 1974), where the value of information is derived endogenously from the agent’s dynamic program, not bolted on as a separate heuristic. The inference tradition also uses Bayesian learning to describe the same phenomenon, but always within a fully specified model of beliefs and preferences; in RL, exploration can be a purely algorithmic device with no decision-theoretic foundation.
@@
Bootstrapping in statistics refers to Efron’s resampling method (Efron, 1979); in RL, it means updating a value estimate using another value estimate rather than a complete realized return, as when a TD algorithm uses the target $r_{t+1}+\gamma V(s_{t+1})$ that depends on the current, uncertain estimate $V(s_{t+1})$ (Sutton, 1988). Function approximation in RL refers to representing value functions or policies using parameterized function classes (linear combinations of basis functions, kernel methods, or neural networks), which statisticians will recognize as sieve estimation or nonparametric series estimation, the approximation of an unknown function by projection onto a finite-dimensional basis. Calibration carries unrelated meanings. In the inference tradition, calibration means choosing model parameters to match a set of empirical moments or stylized facts (Kydland and Prescott 1982). In machine learning, calibration refers to probability calibration, ensuring that predicted probabilities match observed frequencies, a statistical property of a classifier’s outputs.
@@
The term bandit itself carries different mathematical content across disciplines. The classical multi-armed bandit in statistics (Thompson, 1933; Rothschild, 1974; Gittins, 1979) is a Bayesian sequential allocation problem. The state is the agent’s posterior belief over the unknown arm distributions, and the solution is the Gittins index, an optimal allocation rule derived from the theory of optimal stopping. In the RL and computer science literature, bandits are instead framed as frequentist regret-minimization problems. Algorithms such as UCB provide worst-case bounds on cumulative regret $\sum_{t=1}^{T}(\mu^{*}-\mu_{A_{t}})$ without requiring Bayesian priors. The two traditions ask fundamentally different questions, Bayesian optimality of the full sequential problem versus minimax regret rates over adversarial or stochastic environments, and their answers are not directly comparable. Section 10 adopts the regret framework because it connects more naturally to the sample complexity concerns that arise in field experiments.